# Mixed-Strategy Nash Equilibrium for Crowd Navigation

### v2

@@ VIII Conclusion
In this work, we propose a computation-efficient mixed-strategy Nash equilibrium model for crowd navigation. Mixed-strategy Nash equilibrium provides a high-level cooperation model for the robot to plan actions that leverage human cooperation for collision avoidance while maintaining uncertainty in human behavior. Despite the general hardness of computing mixed-strategy Nash equilibrium, we achieve real-time inference speed on a laptop CPU by establishing the formal connection between mixed-strategy Nash equilibrium and a simple iterative Bayesian belief updating scheme. We further enhance the proposed Bayesian updating scheme by extracting temporal Gaussian processes from offline trajectory datasets as part of the Nash equilibrium model, using existing datasets to bring the model closer to actual human behavior. The proposed model can be incorporated into existing navigation frameworks to navigate alongside humans in unstructured environments with static obstacles. Our experiment results show that our model significantly improves the safety of a human-unaware planner without compromising navigation efficiency. Compared to other crowd navigation algorithms, our model consistently outperforms them in both safety and navigation efficiency. Our method also demonstrates human-level crowd navigation performance on top of a meta-planner when interacting with real human crowds in unstructured environments.
@@
Beyond the technical contributions, our work also provides valuable insights into how game theory models can be applied to robotics applications. Most, if not all, existing game-theoretic planning algorithms are designed with a top-down approach: the planner aims to solve a generalized game, with the player objectives being specified later for different applications. While this approach can be applied to arbitrary games, the limitation of the top-down approach is the hardness of solving a generalized game, which is often intractable or too expensive to compute. In this work, we take a bottom-up approach to applying game theory models. We use mixed-strategy Nash equilibrium as a high-level principle to design a specific behavioral model for crowd navigation, with the purpose of supporting real-time decision-making. This bottom-up approach allows us to utilize the analytical power of mixed-strategy Nash equilibrium while maintaining a sufficient computation load for real-time robot navigation. We consider our bottom-up approach makes an orthogonal contribution, with respect to existing top-down approaches, to the principles of designing game-theoretic planners.
@@
There are several limitations of the proposed crowd navigation framework. First, our framework does not support extra constraints on decision-making, such as guarantees on static obstacle avoidance. Even though we have zero static obstacle collision in the human dataset benchmark, this is achieved by using a collision-free meta-planner as the nominal strategy. Our method does not guarantee the robot from a collision with environmental obstacles. Second, the practical performance of our crowd navigation framework is limited by the sampling scheme. We proposed two sampling strategies for approximating the mixed-strategy Nash equilibrium. The rejection sampling-based strategy generates a more accurate posterior but is too slow for real-time decision-making. The importance-sampling-based strategy supports real-time posterior estimation, but its performance is limited by the support of initial samples and suffers from common issues faced by particle filters, such as sample degeneracy. Third, throughout the experiments, we assume the localization for the robot and perception of pedestrians are given, while perception in real human crowds is still an open challenge. Our ongoing work involves testing the algorithm on the hardware in the real world, and investigating how sensor noise would affect the algorithm’s performance.
@@
Mixed-strategy Nash equilibrium paves the way for future works to achieve truly adaptive crowd navigation in varying environments, as it provides extra information enhancing the robot’s adaptivity during the run-time. We want to point out an exciting future direction stemming from our model, which is adaptive or contingent control synthesis based on mixed-strategy Nash equilibrium. In this work, we simply synthesize robot control signals to follow the mean of the converged mixed-strategy of the robot. This approach does not fully utilize the information encoded in mixed-strategy Nash equilibrium. For example, mixed-strategy predictions can serve as a prior distribution when existing information is insufficient for precise behavior prediction. Based on this prior, the robot could quickly refine its prediction or perform contingency planning with new online measurements.
@@
Theorem 2 and Theorem 3 can be used to prove the convergence of Algorithm 1 and Algorithm 2. Again, we start with two-agent social navigation and extend the result to multi-agent crowd navigation.
@@
The sequence of $\{(p_{i}^{[k]},p_{j}^{[k]})\}$ from the iterations of Algorithm 1 is a convergent sequence.
@@
We prove the convergence of the sequence through the monotone convergence theorem, where we show that the sequence monotonically decreases on a function with finite lower bound. We start with the monotone decrease result.
@@
$\mathbb{E}_{p_{i}^{[k+1]},p_{j}^{[k]}}[r]{+}D(p_{i}^{[k+1]}\|p^{\prime}_{i})\leq\mathbb{E}_{p_{i}^{[k]},p_{j}^{[k]}}[r]{+}D(p_{i}^{[k]}\|p^{\prime}_{i})$ (74)
@@
$\mathbb{E}_{p_{i}^{[k+1]},p_{j}^{[k+1]}}[r]{+}D(p_{j}^{[k+1]}\|p^{\prime}_{j})\leq\mathbb{E}_{p_{i}^{[k+1]},p_{j}^{[k]}}[r]{+}D(p_{j}^{[k]}\|p^{\prime}_{j})$ (76)
@@
The above inequality means the sequence $\{(p_{i}^{[k]},p_{j}^{[k]})\}$ from Algorithm 1 monotonically decrease the function:
$R(p_{i}^{[k]},p_{j}^{[k]})+D(p_{i}^{[k]}\|p^{\prime}_{i})+D(p_{j}^{[k]}\|p^{\prime}_{j})$ (79)
Based on Assumption 1 and the non-negativity of KL-divergence, this function (79) has a finite lower-bound. Thus, based on the monotone convergence theorem, the sequence $\{(p_{i}^{[k]},p_{j}^{[k]})\}$ is convergent. ∎∎
Theorem 5 can be extended to the case of multi-agent crowd navigation in Algorithm 2.
@@
The sequence of $\{(p_{0}^{[k]},\dots,p_{N-1}^{[k]})\}$ from the iterations of Algorithm 2 is a convergent sequence.
@@
By iterating the agent index $a$ from 00 to $N-1$, based on the chain rule of inequality, we have the following chain of inequalities:
@@
which means the sequence $\{(p_{0}^{[k]},\dots,p_{N-1}^{[k]})\}$ is monotonically decreases the function:
$\sum_{i=1}^{N}\sum_{j=i+1}^{N}\mathbb{E}_{p_{i}^{[k]},p_{j}^{[k]}}[r]+\sum_{i=1}^{N}D(p_{i}^{[k]}\|p^{\prime}_{i})$ (89)
Based on Assumption 1 and the non-negativity of KL-divergence, this function (89) is lower-bounded. Thus, based on the monotone convergence theorem, the sequence $\{(p_{1}^{[k]},\dots,p_{N}^{[k]})\}$ is convergent. ∎∎
VIII-A Proof for Theorem 1
@@
We prove this by contradiction. Denote the sequence converges to the point $\{(p_{1}^{*},\dots,p_{N}^{*})\}$, assume the convergent point is not a Nash equilibrium, then:
$J_{i}(p_{1}^{*},\dots,p_{i},\dots,p_{N}^{*})<J_{i}(p_{1}^{*},\dots,p_{i}^{*},\dots,p_{N}^{*})$ (90)
$\exists p_{i}\in\mathcal{P},i\in\mathcal{I}$
Based on the definition of the general-sum game player objective (12), the above inequality means $p_{i}^{*}$ is not the global minimum of the following function:
$\mathbb{E}_{p,p_{/i}}[r]+D(p\|p^{\prime}_{i})$ (91)
This contradicts Theorem 3. ∎∎
VIII-B Proof for Theorem 4
@@
Recall in the proof for Theorem 6, the sequence of $\{(p_{0}^{[k]},\dots,p_{N-1}^{[k]})\}$ from the iterations of Algorithm 2 monotonically decreases the function:
$\sum_{i=1}^{N}\sum_{j=i+1}^{N}\mathbb{E}_{p_{i}^{[k]},p_{j}^{[k]}}[r]+\sum_{i=1}^{N}D(p_{i}^{[k]}\|p^{\prime}_{i})$ (92)
@@
$=$ $\sum_{i=1}^{N}\sum_{j=i+1}^{N}R(p^{\prime}_{i},p^{\prime}_{j})+0$ (94)
Appendix B: Complete human dataset benchmark results
@@
Appendix C: Pseudocode
See next page.
Algorithm 3 BRNE navigation (rejection sampling)
2:     $k\leftarrow 0$triangleright $k$ is the negotiation step.
3:     for $i\in[1,N]$ do
4:         for $j\in[1,M]$ do
5:              $\mathbf{s}_{i,j}^{[k]}\leftarrow\mathbf{s}^{\prime}_{i,j}$
6:         end for
7:     end for
8:     while convergence criterion not met do
9:         for $i\in[1,N]$ do
11:              for $j\in[1,M]$ do $\triangleright$ rejection sampling
12:                  $\mathbf{s}_{new}\leftarrow$ Draw new sample from $p^{\prime}_{i}$
13:                  $\omega\leftarrow z(\mathbf{s}_{new}|\mathbf{p}_{/i}^{[k]})$
14:                  $u\leftarrow Uniform(0,1)$
15:                  while $\gamma\cdot u\geq\omega$ do
16:                       $\mathbf{s}_{new}\leftarrow$ Draw new sample from $p^{\prime}_{i}$
17:                       $\omega\leftarrow z(\mathbf{s}_{new}|\mathbf{p}_{/i}^{[k]})$
18:                       $u\leftarrow Uniform(0,1)$
19:                  end while
20:                  $\mathbf{s}_{i,j}^{[k]}\leftarrow\mathbf{s}_{new}$
21:              end for
22:         end for
23:         $k\leftarrow k+1$
24:     end while
25:     return $\mathbf{p}_{1}^{[k]},\dots,\mathbf{p}_{N}^{[k]}$
26:end procedure
Algorithm 4 BRNE navigation (importance sampling)
2:     $k\leftarrow 0$triangleright $k$ is the negotiation step.
3:     for $i\in[1,N]$ do
4:         for $j\in[1,M]$ do
6:         end for
7:     end for
8:     while convergence criterion not met do
9:         for $i\in[1,N]$ do
10:              for $j\in[1,M]$ do $\triangleright$ importance sampling
13:                  $\mathbf{w}_{i,j}^{[k+1]}\leftarrow\frac{1}{N-1}\mathbf{w}$
14:              end for
16:         end for
17:         $k\leftarrow k+1$
18:     end while
19:     return $\mathbf{w}_{1}^{[k]},\dots,\mathbf{w}_{N}^{[k]}$
20:end procedure

### v4

@@ VI Conclusion
In this work, we propose a computation-efficient mixed strategy Nash equilibrium model for crowd navigation. Mixed strategy Nash equilibrium provides a high-level cooperation model for the robot to plan actions that leverage human cooperation for collision avoidance while maintaining the uncertainty in human behavior. Despite the general hardness of computing mixed strategy Nash equilibrium, we achieve real-time inference speed on a laptop CPU and a low-power embedded computer by establishing the formal connection between mixed strategy Nash equilibrium and a simple iterative Bayesian update scheme. We name the proposed model Bayes’ Rule Nash Equilibrium (BRNE). We further develop a model predictive crowd navigation framework using Gaussian processes to bootstrap agents’ nominal mixed strategies. The proposed model can be incorporated into existing navigation frameworks to navigate alongside humans in unstructured environments with static obstacles. Our experiment results show that our BRNE model significantly improves the safety of a human-unaware planner without compromising navigation efficiency in the human dataset benchmark, reaching human-level performance. Compared to other crowd navigation algorithms, our model consistently outperforms them in both safety and navigation efficiency. Lastly, we demonstrate the practicality of our model on an untethered robot quadruped for real-time crowd navigation with fully onboard perception and computation.
@@
Beyond the technical contributions, our work also provides valuable insights into how game theory models can be applied to robotics applications. Most, if not all, existing game-theoretic planning algorithms are designed with a top-down approach: the planner aims to solve a generalized game, with the player objectives being specified later for different applications. While this approach can be applied to arbitrary games, the limitation of the top-down approach is the hardness of solving a generalized game, which is often intractable or too expensive to compute. In this work, we take a bottom-up approach to apply game theory models. We use mixed strategy Nash equilibrium as a high-level principle to design a specific behavioral model for real-time crowd navigation with limited computation resources. This bottom-up approach allows us to utilize the analytical power of mixed strategy Nash equilibrium while maintaining a sufficient computation load for real-time robot navigation. We consider our bottom-up approach to make an orthogonal contribution, with respect to existing top-down approaches, to the principles of designing game-theoretic planners.
@@
There are several limitations of the proposed crowd navigation framework. First, our framework does not support extra constraints on decision-making, such as guarantees on static obstacle avoidance. Even though we have zero static obstacle collision in the human dataset benchmark, this is achieved using a collision-free meta-planner as the nominal strategy. Our method itself does not guarantee the robot from a collision with environmental obstacles. Second, the sampling scheme limits the practical performance of our crowd navigation framework. We proposed two sampling strategies for approximating the mixed strategy Nash equilibrium. The rejection sampling-based strategy generates a more accurate posterior but is too slow for real-time decision-making. The importance-sampling-based strategy supports real-time posterior estimation, but its performance is limited by the support of initial samples and suffers from common issues faced by particle filters, such as sample degeneracy. Third, throughout the experiments, we assume the localization for the robot and perception of pedestrians are given, while perception in real human crowds is still an open challenge. Our ongoing work involves benchmarking the algorithm on the hardware in the real world, and investigating how sensor noise would affect the algorithm’s performance.
@@
Mixed strategy Nash equilibrium paves the way for future works to achieve truly adaptive crowd navigation in varying environments, as it provides extra information enhancing the robot’s adaptivity during the run-time. We want to point out an exciting future direction stemming from our model, which is adaptive or contingent control synthesis based on mixed strategy Nash equilibrium. In this work, we simply synthesize robot control signals to follow the mean of the converged mixed strategy of the robot. This approach does not fully utilize the information encoded in mixed strategy Nash equilibrium. For example, mixed strategy predictions can be a prior distribution when existing information is insufficient for precise behavior prediction. Based on this prior, the robot could quickly refine its prediction or perform contingency planning with new online measurements.
Acknowledgements
The authors would like to thank Matthew Elwin and Maia Traub for their assistance with the hardware demonstration.
@@ Acknowledgements
Theorem 2 and Theorem 3 are to prove the convergence of Algorithm 1 and Algorithm 2. Again, we start with the two-agent scenario and extend the result to the multi-agent scenario.
@@
The sequence of $\{(p_{i}^{[k]},p_{j}^{[k]})\}_{k}$ from the iterations of Algorithm 1 is a convergent sequence.
@@
We prove the sequence’s convergence through the monotone convergence theorem, by proving that the sequence monotonically decreases on a function with a finite lower bound. We start with the monotone decrease result.
@@
$\mathbb{E}_{p_{i}^{[k+1]},p_{j}^{[k]}}[r]{+}D(p_{i}^{[k+1]}\|p^{\prime}_{i})\leq\mathbb{E}_{p_{i}^{[k]},p_{j}^{[k]}}[r]{+}D(p_{i}^{[k]}\|p^{\prime}_{i}).$ (33)
@@
$\mathbb{E}_{p_{i}^{[k+1]},p_{j}^{[k+1]}}[r]{+}D(p_{j}^{[k+1]}\|p^{\prime}_{j})\leq\mathbb{E}_{p_{i}^{[k+1]},p_{j}^{[k]}}[r]{+}D(p_{j}^{[k]}\|p^{\prime}_{j}).$ (35)
@@
The above inequality means the sequence $\{(p_{i}^{[k]},p_{j}^{[k]})\}_{k}$ from Algorithm 1 monotonically decreases the function:
$F(p_{i},p_{j})=\mathbb{E}_{p_{i},p_{j}}[r]+D(p_{i}\|p^{\prime}_{i})+D(p_{j}\|p^{\prime}_{j}).$ (38)
Based on the non-negativity of KL-divergence, this function (38) has a finite lower-bound. Thus, based on the monotone convergence theorem, the sequence $\{(p_{i}^{[k]},p_{j}^{[k]})\}_{k}$ is convergent under (20). ∎∎
Theorem 5 can be extended to the multi-agent scenario in Algorithm 2.
@@
The sequence of $\{(p_{1}^{[k]},\dots,p_{N}^{[k]})\}_{k}$ from the iterations of Algorithm 2 is a convergent sequence.
@@
$+{\sum_{i=a+1}^{N}\sum_{j=a}^{N-1}}\mathbb{E}_{p_{i}^{[k]},p_{j}^{[k]}}[r]$
$+\sum_{i=1}^{a-1}D(p_{i}^{[k+1]}\|p^{\prime}_{i})+\sum_{j=i+1}^{N}D(p_{i}^{[k]}\|p^{\prime}_{i}).$ (45)
By iterating the agent index $a$ from $1$ to $N$, based on the chain rule of inequality, we have the following chain of inequalities:
@@
which means the sequence $\{(p_{0}^{[k]},\dots,p_{N-1}^{[k]})\}_{k}$ monotonically decreases the function:
$F(p_{1},\dots,p_{N})=\sum_{i=1}^{N}\sum_{j=i+1}^{N}\mathbb{E}_{p_{i},p_{j}}[r]+\sum_{i=1}^{N}D(p_{i}\|p^{\prime}_{i}).$ (48)
Based on the non-negativity of KL-divergence, this function (48) is lower-bounded. Thus, based on the monotone convergence theorem, the sequence $\{(p_{1}^{[k]},\dots,p_{N}^{[k]})\}_{k}$ is convergent under (20). ∎∎
VI-A Proof for Theorem 1
@@
Theorem 6 has proved the convergence of the mixed strategy sequence under (20). By contradiction, we can prove the limit point is a global Nash equilibrium of the mixed strategy game (18). Denote the sequence converges to the limit point $(p_{1}^{*},\dots,p_{N}^{*})$, assume the limit point is not a global Nash equilibrium of (18), then
[…]
$\mathbb{E}_{p_{i},p_{/i}^{*}}[r]+D(p_{i}\|p^{\prime}_{i})<\mathbb{E}_{p_{i}^{*},p_{/i}^{*}}[r]+D(p_{i}^{*}\|p^{\prime}_{i}).$ (49)
Since the mixed strategy sequence from Algorithm 2 monotically decreases the lower-bounded function (48) and the right hand side of (49) is part of the summation in (48), the inequality (49) indicates $(p_{1}^{*},\dots,p_{N}^{*})$ is not a limit point under (20), which contradicts the assumption, thus completes the proof. ∎∎
VI-B Proof for Theorem 4
@@
Recall in the proof for Theorem 6, the sequence of $\{(p_{1}^{[k]},\dots,p_{N}^{[k]})\}$ from the iterations of Algorithm 2 monotonically decreases the function:
$\sum_{i=1}^{N}\sum_{j=i+1}^{N}\mathbb{E}_{p_{i}^{[k]},p_{j}^{[k]}}[r]+\sum_{i=1}^{N}D(p_{i}^{[k]}\|p^{\prime}_{i}).$ (50)
@@
$=$ $\sum_{i=1}^{N}\sum_{j=i+1}^{N}\mathbb{E}_{p^{\prime}_{i},p^{\prime}_{j}}[r]+0,$ (52)
which completes the proof. ∎∎
Appendix B: Pseudocode
Algorithm 3 BRNE navigation (rejection sampling)
2:     $k\leftarrow 0$triangleright $k$ is the negotiation step.
3:     for $i\in[1,N]$ do $\triangleright$ Number of agents
4:         for $j\in[1,M]$ do $\triangleright$ Number of samples///agent
5:              $\mathbf{s}_{i,j}^{[k]}\leftarrow\mathbf{s}^{\prime}_{i,j}$triangleright Sample from nominal strategy
6:         end for
7:     end for
8:     while convergence criterion not met do
9:         for $i\in[1,N]$ do
11:              for $j\in[1,M]$ do $\triangleright$ Rejection sampling
12:                  $\mathbf{s}_{new}\leftarrow$ Draw new sample from $p^{\prime}_{i}$
13:                  $\omega\leftarrow z(\mathbf{s}_{new}|\mathbf{p}_{/i}^{[k]})$
14:                  $u\leftarrow Uniform(0,1)$
15:                  while $\gamma\cdot u\geq\omega$ do
16:                       $\mathbf{s}_{new}\leftarrow$ Draw new sample from $p^{\prime}_{i}$
17:                       $\omega\leftarrow z(\mathbf{s}_{new}|\mathbf{p}_{/i}^{[k]})$
18:                       $u\leftarrow Uniform(0,1)$
19:                  end while
20:                  $\mathbf{s}_{i,j}^{[k]}\leftarrow\mathbf{s}_{new}$
21:              end for
22:         end for
23:         $k\leftarrow k+1$
24:     end while
25:     return $\mathbf{p}_{1}^{[k]},\dots,\mathbf{p}_{N}^{[k]}$
26:end procedure
Algorithm 4 BRNE navigation (importance sampling)
2:     $k\leftarrow 0$triangleright $k$ is the negotiation step.
3:     for $i\in[1,N]$ do $\triangleright$ Number of agents
4:         for $j\in[1,M]$ do $\triangleright$ Number of samples///agent
6:         end for
7:     end for
8:     while convergence criterion not met do
9:         for $i\in[1,N]$ do
10:              for $j\in[1,M]$ do $\triangleright$ Importance sampling
13:                  $\mathbf{w}_{i,j}^{[k+1]}\leftarrow\mathbf{w}$
14:              end for
16:         end for
17:         $k\leftarrow k+1$
18:     end while
19:     return $\mathbf{w}_{1}^{[k]},\dots,\mathbf{w}_{N}^{[k]}$
20:end procedure
@@
Appendix C: Complete human dataset benchmark results